two cars, 645 kilometers apart, begin driving toward each other on long straight highway. one car travels 100 kilometers per hour and the other 10 kilometers per hour. At the same time, a canary, starting on one car, flies back and forth between the two cars as they approach each other. If the canary flies 120 kilometers per hour and spends no time to turn around at each car, how far has it flown when the cars collide?
Answer by mananth(16949)

We have to find out how much time the Canary flies, before they collide
Car I speed = 100 km/h
Car II speed = 10km/h
they are moving towards each other
so add up their speeds= 110 km/h
distance = 645 km
Time = d/r
t= 645/110
t= 5.86 hours
speed of Canary = 120 km/h
time = 5.86 hours
so distance it flies = 5.86 *120= 703 km approx